1.In an enter - baranggay basketball league,the team from baranggay Carolina has won 15 out of 25 games,a winning percentage of 60%. We have seen that they need to win 25 more games consecutively to raise their percentage for at least 80%.


a.what function will represent the winning percentage of baranggay Carolina if x is the number of remaining games?


b.what will be the winning persentage after 15?20?25?30?35 games?


c.plot the ordered pairs you got in questions (a)and(b) in a Cartesian plane and connect it using a smooth curve


1
Expert's answer
2021-11-22T18:12:22-0500


a.

if x is the number of remaining games

number of games played = 50 - x

number of games won = 40 - x 

function for winning percentage

"f(x)=\frac{number\space of\space games \space won}{number \space of \space games\space played}\times 100"


"=\frac{40-x}{50-x}\times100"


b.

after 15 game , remaining game = x = 50 - 15 = 35

"=\frac{40-x}{50-x}\times100\\=\frac{40-35}{50-35}\times100\\=\frac{100}{3}\%"

after 20 games . x = 50 - 20 = 30

"=\frac{40-x}{50-x}\times100\\=\frac{40-30}{50-30}\times100\\=50\%"

after 25 games , x = 50 - 25

"f(25)=\frac{40-25}{50-25}\times100\\=60\%"

after 30 games , x = 50-30=20

"f(20)=\frac{40-20}{50-20}\times100\\=\frac{200\%}{3}"

after 35 games x = 50 - 35 = 15

"f(15)=\frac{40-15}{50-15}\times100\\=\frac{500\%}{7}"